OpenStudy (anonymous):

the results of a medical test show that of 32 people selected at random who were given the test, 2 tested positive and 30 tested negative. Determine the odds in favor of a person selected at random testing positive on the test.

OpenStudy (kropot72):

The odds are the ratio of the probability of an event occurring to that of its not occurring. What is the experimental probability of a randomly tested person testing positive?

OpenStudy (anonymous):

you lost me with that last part.

OpenStudy (kropot72):

The experimental probability of a randomly tested person testing positive is given by: \[\large \frac{number\ testing\ positive}{total\ number\ tested}\]

OpenStudy (anonymous):

so it would be as simple as 2/64?

OpenStudy (anonymous):

then simplify it obviously

OpenStudy (kropot72):

Where did '64' come from?

OpenStudy (anonymous):

crap so it would just be "32" sorry insomnia is setting in!

OpenStudy (kropot72):

Yes, the experimental probability of a randomly tested person testing positive is 2/32. Next step: What is experimental probability of a randomly tested person testing negative?

OpenStudy (anonymous):

so it would be 30/32?

OpenStudy (anonymous):

testing negative.

OpenStudy (kropot72):

Correct. So looking at the definition of odds: 'The odds are the ratio of the probability of an event occurring to that of its not occurring.' So an initial result for the required odds in favor of a person selected at random testing positive on the test is: 2/32 : 30/32 which can be simplified. Can you simplify it?

OpenStudy (kropot72):

The aim is to simplify \[\large \frac{2}{32}:\frac{30}{32}\] to get an integer on each side.

OpenStudy (kropot72):

Multiply each term by 32/2

OpenStudy (kropot72):

\[\large (\frac{2}{32}\times\frac{32}{2}):(\frac{30}{32}\times\frac{32}{2})=?\]

OpenStudy (anonymous):

Sorry I was reading the book and it gave me a weird formula I was trying to wrap my head around based on what we were working on

OpenStudy (anonymous):

it should be 1:15 if I did my math right

OpenStudy (kropot72):

Correct :)

OpenStudy (anonymous):

Theres a formula for odds in favor, the way you just walked out, is that the same process?

OpenStudy (kropot72):

If the probability of an event A occurring is P(A) and the probability of event A not occurring is \[\large P(\bar{A})\] then the odds in favor of event A is given by \[\large P(A):P(\bar{A})\] This is the method that I used.

OpenStudy (anonymous):

awesome, ok I think I have it. up for helping me with a couple more?

OpenStudy (kropot72):

You're welcome :)